What is the probability that a corporate tax return will be audited?
The answer will depend on the tax jurisdiction. Different countries have different resources allocated for this purpose.
The answer will depend on the tax jurisdiction. Different countries have different resources allocated for this purpose.
The answer will depend on the tax jurisdiction. Different countries have different resources allocated for this purpose.
The answer will depend on the tax jurisdiction. Different countries have different resources allocated for this purpose.
What is the probability of selecting a king or queen?
It depends on the context which is not specified. If picking a topic for a history project on the monarchy of some country it would be 1!
If it concerns selecting one card, at random, from a standard deck of playing cards, the answer is 8/52 = 2/13.
The increase in sample size will reduce the confidence interval.
The increase in standard deviation will increase the confidence interval.
The confidence interval is not based on a linear function so the overall effect will require some calculations based on the levels before and after these changes.
It would depend on the relative rates at which the change in sample size and change in standard deviation occurred. If the sample size increased more quickly than then standard deviation, in some sense, then the size of the confidence interval would decrease. Conversely, if the standard deviation increased more quickly than the sample size, in some sense, then the size of the confidence interval would increase.
What is the probability of drawing a red ten or a black ace?
The answer depends on
Thus, if 49 cards are drawn without replacement from an ordinary deck, whether randomly or not, the probability is 1.
For a single card drawn randomly, the probability is 1/13.
The answer depends on
Thus, if 49 cards are drawn without replacement from an ordinary deck, whether randomly or not, the probability is 1.
For a single card drawn randomly, the probability is 1/13.
The answer depends on
Thus, if 49 cards are drawn without replacement from an ordinary deck, whether randomly or not, the probability is 1.
For a single card drawn randomly, the probability is 1/13.
The answer depends on
Thus, if 49 cards are drawn without replacement from an ordinary deck, whether randomly or not, the probability is 1.
For a single card drawn randomly, the probability is 1/13.
How many different ways can 360 players in a marching band be arranged in a rectangle?
If you consider an 18 x 20 rectangle to be the same as a 20 x 18, there are twelve ways. If not, double that.
What is the probibility of 1 mean?
A probability of 1 refers to an event that is certain to happen.
A probability of 1 refers to an event that is certain to happen.
A probability of 1 refers to an event that is certain to happen.
A probability of 1 refers to an event that is certain to happen.
What is the probability of flipping a head when you roll a coin 3 times?
If it is a fair coin, the probability of getting at least one Head from 3 flips is 7/8
If it is a fair coin, the probability of getting at least one Head from 3 flips is 7/8
If it is a fair coin, the probability of getting at least one Head from 3 flips is 7/8
If it is a fair coin, the probability of getting at least one Head from 3 flips is 7/8
What is the probability of getting one head when flipping a coin three times?
The probability is 3/8.
The probability is 3/8.
The probability is 3/8.
The probability is 3/8.
What is probability of getting a 3 in a deck of cards?
For an ordinary deck of cards, the probability is 1. All decks of playing cards contain 3s.
For an ordinary deck of cards, the probability is 1. All decks of playing cards contain 3s.
For an ordinary deck of cards, the probability is 1. All decks of playing cards contain 3s.
For an ordinary deck of cards, the probability is 1. All decks of playing cards contain 3s.
What is the probability that the number rolled will be greater that 4?
The probability that the number rolled, on a fair, six sided die, will be greater than 4 is 1/3.
The probability that the number rolled, on a fair, six sided die, will be greater than 4 is 1/3.
The probability that the number rolled, on a fair, six sided die, will be greater than 4 is 1/3.
The probability that the number rolled, on a fair, six sided die, will be greater than 4 is 1/3.
6
2 on the coin, 3 on the pointer: 2x3=6
What is the probability of getting exactly three heads when you flip a coin four times?
If it is a fair coin, the probability is 1/4.
If it is a fair coin, the probability is 1/4.
If it is a fair coin, the probability is 1/4.
If it is a fair coin, the probability is 1/4.
The probability of (1 or 2 or 3) on the first (or any) roll is 1/2 = 50% .
The probability of (4 or 5 or 6) on the second (or any) roll is 1/2 = 50% .
The probability of exactly that result on two rolls is (1/2 x 1/2)= (50% x 50%) = 1/4 = 25% .
How many different ways can 4 people be seated on a row of seats?
4 people seated on a row may have 4! = 4x3x2x1 = 24 different ways to sit.
How are experimental probability and theoretical probability alike?
They are both measures of how likely it is that a particular event will occur.
They are both measures of how likely it is that a particular event will occur.
They are both measures of how likely it is that a particular event will occur.
They are both measures of how likely it is that a particular event will occur.
What is the probability of winning a lotto?
The answer depends on the lotto. The relevant variables are:
What is the probability of getting 4 of a kind if 5 identical dice are tossed simulataneously?
If they are fair dice, the probability is 0.0032
If they are fair dice, the probability is 0.0032
If they are fair dice, the probability is 0.0032
If they are fair dice, the probability is 0.0032
What is the probability to roll a 5?
The answer depends on how many dice are rolled, how often and whether they are fair dice.
For a single roll of a fair die the answer is 1/6.
The answer depends on how many dice are rolled, how often and whether they are fair dice.
For a single roll of a fair die the answer is 1/6.
The answer depends on how many dice are rolled, how often and whether they are fair dice.
For a single roll of a fair die the answer is 1/6.
The answer depends on how many dice are rolled, how often and whether they are fair dice.
For a single roll of a fair die the answer is 1/6.
What are the properties of z in a normal distribution?
Z is a variable with mean 0 and variance 1.
Z is a variable with mean 0 and variance 1.
Z is a variable with mean 0 and variance 1.
Z is a variable with mean 0 and variance 1.
How do you find probabilty of negative z score?
Let z be positive so that -z is the negative z score for which you want the probability.
Pr(Z < -z) = Pr(Z > z) = 1 - Pr(Z < z).